OpenStudy (anonymous):

Assessment Help, i will medal First question; Which fraction is equivalent to 1 5/9? A. 41/9 B. 75/3 C. 75/9 D. 68/9

OpenStudy (anonymous):

@Danisaur

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

To rewrite your fraction in this form: \[\frac{ a }{ b }\] You must first, multiply the denominator to its preceding whole: \[1\frac{ 5 }{ 9 } ~ \rightarrow ~ 9 \times 1 = 9\] Then, we add this value to the initial numerator: \[9 + 5 = 14\] This will be our rewritten numerator. We just place this on top of the initial denominator and get our answer: \[\frac{ 14 }{ 9 }\] This isn't one of your choices which suggest that you must have written a typo somewhere @MissNoName.

OpenStudy (anonymous):

no i have not, in my lesson it clearly says 41/9 not 14/9

OpenStudy (anonymous):

Ah, you copied the prompt wrong. You said 1 5/9 when it should have been 7 5/9. Just follow the process I've explained, replace 1 with 7, and you'll get your answer. Please, try to avoid typos in the future. It hinders the both you and I.

OpenStudy (anonymous):

So im thinking A?